OpenStudy (anonymous):

please help......find the plane that passes through 2.0,1 and is parallel to the plane 2x-4y-5z+8 =0 I worked out the vectors (2 , -4 , -5 ) ? what is the next step ?

OpenStudy (anonymous):

any plane parallel to the given plane is 2x-4y-5z+c=0 Because it passes through (2,0,1) 2*2-4*0-5*1+c=0 4-5+c=0 c=1 required eq. is 2x-4y-5z+1=0

OpenStudy (anonymous):

thanks a lot makes sense so you actuallyy are looking for the equation

OpenStudy (anonymous):

when eq is parallel then only constant will change ,direction cosines of normal will remain the same.